# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

### v2

@@ Abstract
To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity. These techniques are often designed with a pre-defined KV budget; however, as the optimal budget varies by different input lengths and task types, the existence of a fixed budget could result in inconsistent performance accepting inputs of diverse domains. To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process. Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average. Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
@@ 1 Introduction
Most large language models (LLMs) follows autoregressive generation, where they rely on KV Cache to store intermediate states during inference [1, 2, 3]. Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache [4, 5, 6]. However, when the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally. Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference. For instance, Llama3 8B model [7] requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
Consequently, recent works have proposed to effectively reduce KV cache, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary. These observations have led to several optimization methods, such as H2O [8], ScissorHands [9], StreamingLLM [10], and FastGen [11], where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria. Apart from pruning, other paradigms have been proposed as well to reduce KV consumption, such as KVMerger [12] and D2O [13] that merge or compress KV vectors instead of hard-pruning.
While these studies have successfully demonstrated that a decent portion of KV cache can be discarded in practice, there remains one common issue that has yet to be addressed: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs. In real-world scenarios, one would need to carefully tune budget thresholds optimized for different domains; otherwise, setting a uniform threshold in diverse domains may potentially bring large degradation on certain tasks. We reckon the need for a pre-defined threshold as an overlooked limitation that can lead to unstable and inconsistent performance in practice.
Table 1: With pre-defined cache budgets, KV pruning methods on Llama3 exhibit inconsistent performance across domains (percentage relative to full-cache scores).
Methods GSM8K GPQA CoQA
@@
H2O 3.2% 8.5% 96.5%
SLM 4.3% 11.5% 76.1%
SnapKV 1.7% 3.9% 98.6%
To further illustrate, Table 1 shows that when retaining 90% KV cache, Llama3-8B-Instruct could achieve near full-cache performance across three domains. However, the performance is shown inconsistent with 20% budget; especially on the math benchmark GSM8K [14], all three KV pruning methods drastically fall to less than 5% relative to the full-cache performance.
To address the aforementioned limitation, in this work, we propose a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget. In particular, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation. A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
Our method features a two-step process: (1) after the prefilling stage, rank all input tokens according to an importance metric; (2) discard the KV cache of each token sequentially, until hitting a stopping criterion. In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle [10, 15]; ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation [1, 2, 8].
To eliminate the pre-defined KV budget while always targeting full-cache performance, the stopping criterion becomes crucial, such that it should signal whenever the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling the objective to approximate full performance with dynamic cache pruning. In DBudgetKV, we design a simple attention-based metric as the stopping condition: the metric compares the difference between the Frobenius form of the full attention matrix and the reduced attention matrix of which pruned positions’ scores are set to zero. When the difference exceeds a universal threshold, the pruning process halts. In practice, we find that this metric empirically correlates well with the full-cache performance regardless of input types, serving as an estimator to bridge the pruning state and the subsequent generation.
@@
Figure 1: The overall workflow of DBudgetKV in Section 3.2. After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved to subsequent generation.
To examine the performance consistency across diverse domains, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g. mathematical and commonsense reasoning, reading comprehension and coding, demonstrating that DBudgetKV achieves our objective effectively and robustly. With dynamic KV budgets, the resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 [7], Qwen2.5 [16], and Mistral [17]. The dynamic pruning of DBudgetKV automatically allocates relatively high cache budget on math benchmarks to prevent degradation, while achieving high compression ratio on comprehension tasks by up to 85%. With Llama3-8B, DBudgetKV reaches an average budget ratio of 63.7% with minimal degradation across inputs. Overall, empirical results suggest that our proposed method is able to accomplish lossless KV cache pruning with strong generalizability. Besides, it does not rely on specific model architectures, and is orthogonal to other KV optimization techniques. We further conduct experiments for in-depth analysis and ablation studies, where we validate that each step of DBudgetKV is necessary in achieving the desired performance, depicting the rationality of various design choices and hyper-parameters. Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to call for the objective that ensures full-cache performance while maximizing the KV cache pruning, addressing the limitation of inconsistent performance with fixed KV budget thresholds.
  • We introduce DBudgetKV that employs a straightforward two-step process, where the dynamic pruning is halted with minimal degradation to the generation performance.
  • Extensive evaluation with diverse tasks and models indicates that compared to prior works, our method achieves lossless KV cache compression effectively, efficiently and robustly.
2 Related Work
KV Cache Compression
Recent advancements in KV cache compression have emerged to address memory constraints for model inference. Scissorhands [9] observes the repetitive nature of attention patterns and the persistence of token importance. It then preserves the most important tokens based on attention scores to achieve KV cache compression. H2O [8] introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters". StreamingLLM [10] enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens. SnapKV [18] enhances compression efficiency through attention score-based selection and clustering of critical KV positions. FastGen [11] proposes an adaptive KV cache management system that customizes retention strategies per attention head. Other paradigms beyond KV pruning have been proposed along the way, such as KV cache merging [19].
As prior KV pruning methods typically require setting a pre-defined memory budget, they are by design inherently unable to ensure consistent performance across various domains and datasets, which could prevent their widespread deployment in practice. Our method targets to lift such constraint, achieving near full-cache performance and dynamic budget size without the need of additional tuning.
Attention Pattern
The study of attention head specialization has paved the way for optimizing memory and computation in Transformer models. Early research on encoder architectures shows that some attention heads have persistent positional biases, like consistently focusing on adjacent tokens, indicating redundancy in their learned patterns. Rho-1 [20] reveals that the importance of individual heads varies significantly across tasks, suggesting that not all attention computations are equally crucial during inference, and reveal the potential for dynamically pruning or approximating redundant computations. Recent works [6, 21] have shifted the focus to decoder-only models, where the quadratic memory overhead of retaining intermediate KV vectors poses a significant challenge. Focused Transformer [22] extends the context length of LLMs through a contrastive learning-inspired training process. Landmark-Attention [23] uses a “landmark” token to represent each block of the input and trains the attention to use it for selecting relevant blocks, enabling retrieval of blocks directly through the attention mechanism instead of by relying on a separate mechanism.
Our work leverages attention’s structural regularities, particularly the predictable, position-specific behaviors of certain heads, to dynamically reduce the memory footprint of cached KV states. By identifying and exploiting the properties with position bias and attention distribution, we enable efficient inference without compromising model performance, aligning with the broader goal of dynamic KV cache pruning for various inputs.
@@ 3 Methodology
We first elaborate our objective in Sec. 3.1 as a new direction for KV cache compression. Distinct from previous works, we aim to compress the KV cache as much as possible while ensuring LLMs’ performance remains intact. We then delineate our proposed approach along with its motivation behind in Sec. 3.2. Important implementation details are next discussed in Sec. 3.3.
3.1 Objective of KV Cache Compression
As mentioned in Sec. 1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly [24, 18], or retaining positions by a fixed ratio of input length [8, 13]. While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
It is forthcoming that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions. For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget. Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget.
As a side effect, inference with fixed KV cache pruning may experience performance instability across different inputs, which can be observed by our analysis in Sec. 4.2. We regard it as a practical limitation that motivates the inception of this work.
In addressing this limitation, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget. We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance. In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyper-parameters, which we hope would further advance the research development of KV cache compression techniques.
3.2 DBudgetKV
Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, comprising two steps. First, all tokens are ranked according to an importance metric per Transformers layer. Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion. Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache; rather, it relies on the stopping criterion to determine dynamically, which should work uniformly regardless of the inputs.
The motivation behind our two-step procedure comes from two phenomena brought by previous works. First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM [10]. Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques [9, 8, 18]. For our objective, we propose to combine the two important properties in our approach, described in details as follows.
Importance Ranking
Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head. As the pruning process starts from the least important KV vectors, all tokens are firstly ranked according to an importance strategy.
To this end, we resort to a simple ranking strategy leveraging the position bias, without undergoing any additional computation. As identified by StreamingLLM, a significant portion of attentions is allocated to the initial tokens, regardless of their relevance to the language modeling task, known as attention sinks. Inspired by this, we employ a purely position-based importance evaluation strategy: we assume that the first $m$ tokens are always crucial; while among the remaining $n-m$ tokens, those towards the end are more significant. By this strategy, the importance in $X$ is ranked from the highest to the lowest as $\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$, denoted as $\widehat{X}$, where $m$ is a chosen hyperparameter that works regardless of specific input sequences.
Like previous KV compression works, our choice on importance ranking is empirically supported rather than theoretically based. We have also experimented other attention-based importance ranking, which are provided in Sec. 4.5. The choice of position-based ranking is empirically shown both empirically effective and superior on computational overhead.
Stopping Criterion
With the initially ranked KV vectors, our method then sequentially decides whether to continue or terminate pruning, which requires an evaluation metric as the stopping criterion. Since one evidently cannot know the precise impact of a KV cache position before the generation, we adopt a metric that empirically correlates well with the performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full cache performance.
Inspired by Devoto et al. [25], we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically. For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
We conduct preliminary experiments to validate its reliability. We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero. As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains. The simple Frobenius norm works empirically well in subsequent large-scale evaluation in Sec. 4.

### v3

# Towards Threshold-Free KV Cache Pruning

@@ Abstract
To reduce memory consumption during LLM inference, prior works have proposed numerous methods that focus on KV cache pruning based on various criteria. While these techniques often accomplish lossless memory reduction on many datasets, they often rely on an under-emphasized condition: a dataset/domain-specific budget size threshold needs to be pre-determined to achieve the optimal performance. However, such input-specific tuning may be considerably limited in real-world scenarios, as open-domain inputs span diverse domains, lengths and difficulty levels, without clear boundaries for pre-tuning. Thus, the dependence of an input-sensitive threshold can be an inherent limitation that may cause large degradation on arbitrary inputs.
In this work, we propose a new objective that lifts the threshold constraints for robust KV pruning, calling for “threshold-free” methods that automatically adjust budget sizes while ensuring full-cache performance. We then propose a novel method ReFreeKV as the first solution fulfilling this objective, validated by intensive experiments on 13 datasets of diverse context lengths, task types, and model sizes.
Towards Threshold-Free KV Cache Pruning
Xuanfan Nithanks:  Equal contribution. Partial work done during Xuanfan’s internship at WeChat AI.   Liyan Xu11footnotemark: 1333 Project lead: Liyan Xu <liyanlxu@tencent.com>    Chenyang Lyu   Longyue Wang Mo Yu   Lemao Liu   Fandong Meng   Jie Zhou   Piji Li Nanjing University of Aeronautics and Astronautics WeChat AI, Tencent     Independent Researcher xuanfanni@gmail.com   liyanlxu@tencent.com
@@ 1 Introduction
As current Transformer-based large language models (LLMs) employs autoregressive generation, they produce KV Cache to store intermediate states during inference DBLP:journals/corr/abs-2307-03109; DBLP:journals/corr/abs-2303-08774; DBLP:journals/corr/abs-2402-06196, where they access the cached key and value vectors of past tokens that typically reside within GPU memory for attention calculation DBLP:conf/nips/VaswaniSPUJGKP17; DBLP:journals/corr/abs-1911-02150; DBLP:conf/emnlp/AinslieLJZLS23. Consequently, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference, as they can become massive and grow proportionally when the model size and the input length increase. For instance, Llama3-8B model DBLP:journals/corr/abs-2407-21783 requires 1GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
Focusing on this research topic, numerous methods have been recently proposed to effectively reduce the KV footprint after LLM prefilling. Leveraging the sparsity of attention, retaining full KV cache is not always necessary. Several optimization methods, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23, ScissorHands DBLP:conf/nips/LiuDLWXXKS23, SnapKV DBLP:conf/nips/LiHYVLYCLC24, FastGen DBLP:conf/iclr/Ge0LZ0024, CAKE DBLP:conf/iclr/QinCLHFCLL25, etc., discard the less critical cache positions based on certain pruning criteria. Other paradigms such as KVMerger DBLP:journals/corr/abs-2407-08454 and D2O DBLP:journals/corr/abs-2406-13035 resort to merge or compress KV vectors instead of hard-pruning for achieving the memory reduction effects.
While these studies have confirmed on a number of datasets that a decent portion of KV cache can be pruned, they hinge on a fundamental yet often under-emphasized condition: a threshold, often being dataset-specific, is typically involved for selective tuning to achieve satisfactory results. For instance, D2O takes a pre-defined KV cache budget ratio of 20% to reach full cache performance on LongBench DBLP:conf/acl/BaiLZL0HDLZHDTL24, whereas the required budget jumps to 80% on GSM8K DBLP:journals/corr/abs-2110-14168; likewise, a pre-selected 1024 cache positions is enough for CAKE on LongBench, yet the same budget underperforms on the needle-in-the-haystack test needle.
The existence of such threshold serves fine in idealized settings given dedicated datasets. However, its applicability may be considerably limited in real-world scenarios, where the inputs are intermixed across different domains, lengths and difficulty levels without explicit separation. As a result, optimal thresholds cannot be pre-determined for diverse real-world inputs, potentially making the system less robust and prone to significant performance degradation. For demonstration, Table 1 presents preliminary experiments using H2O, SnapKV and CAKE where inputs from different datasets are mixed to preclude dataset-specific threshold tuning. By varying the KV cache budget ratios, the performance across datasets is shown inconsistent, with a trivial gap under 90% budget but drastically escalating under 20% budget.
Methods GSM8K GPQA CoQA AVG.
@@
H2O 3.2% 8.5% 96.5% 36.1%
SnapKV 1.7% 3.9% 98.6% 34.7%
CAKE 2.2% 15.4% 96.1% 37.9%
Table 1: KV pruning methods that depend on a preset KV budget threshold can exhibit inconsistent performance across domains (percentage relative to full-cache scores using Llama3), making input-specific threshold selection unavoidable for achieving optimal inference.
In this work, we propose a new objective that lifts the threshold dependency in KV cache pruning, thereby enabling the system to robustly handle arbitrary inputs without input-specific tuning. Specifically, our objective prioritizes two principles: 1) the method operates with a universal threshold insensitive to inputs, effectively rendering it “threshold-free”; and 2) the method should consistently achieve performance comparable to its full-cache counterpart, by automatically adjusting KV cache budgets. Only after satisfying these two criteria should the method pursue the best possible compression ratio. As shown by our full experimental results in Table 2, while prior methods may achieve strong compression on certain datasets, none satisfies both critical aspects by our new objective.
Towards our goal, we then propose a novel method, ReFreeKV, featuring a thReshold-Free KV cache pruning. Our method adopts a two-step process with a novel metric to dynamically control the KV cache budget. Concretely, it first ranks all KV cache positions based on their positional importance, a signal proven useful by StreamingLLM DBLP:conf/iclr/XiaoTCHL24; then, it progressively retains key-value vectors in order, and discards the remaining KV cache once the stopping criterion of our proposed input-insensitive Uni-Metric. For minimal overhead, ReFreeKV is implemented with Torch parallel operators, and our analysis in Section 3.3 shows that its latency is on par with prior efficient KV pruning methods varying batch sizes.
The “threshold-free” aspect stems from the empirical validation of our Uni-Metric described in Section 2.2, whose design is shown to be consistent and insensitive to variations in input domains and sequence lengths. Specifically, it adopts a Universal threshold leveraging the Frobenius norm of attention matrix, such that halting by the same threshold empirically incurs minimal degradation relative to the full-cache capacity. By design, ReFreeKV naturally yields higher compression ratios for simpler tasks, while allocating more cache resources to more complex ones.
@@
Figure 1: The overall workflow of ReFreeKV in Section $2.$ After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved to subsequent generation.
To examine its efficacy, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g. mathematical, commonsense reasoning, reading comprehension and coding, demonstrating that ReFreeKV fulfills our objective: without an input-specific threshold, the resulting inference is highly comparable or even surpasses the full-cache performance, evaluated with multiple LLMs of different sizes. Using Llama3-8B, it automatically allocates an average KV budget ratio of 63.7%, while retaining 100% full-cache performance across 13 datasets. Extensive experiments and ablation studies empirically demonstrate that ReFreeKV effectively accomplishes our proposed objective. Moreover, as ReFreeKV does not rely on specific model architectures, it has the potentials to further combine with other orthogonal memory optimization paradigms such as quantization DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2502-04420.
@@ 2 Methodology
In this section, we first elaborate our motivation and the new objective for KV cache compression that differs from prior works. We then delineate our proposed approach ReFreeKV, along with its key implementation details.
2.1 Threshold-Free: A New Objective
As aforementioned, previous KV cache compression methods require a pre-selected budget threshold. While these methods can perform well on numerous datasets, it is forthcoming that the dependence on a threshold can become an inherent limitation. As the optimal threshold can vary across different inputs, it may not be feasible for such systems to pick appropriate thresholds or maintain stable performance in practical scenarios involving arbitrary inputs and open-domain instructions. Certain inputs may necessitate a relatively higher memory budget, such as tasks with multi-step or mathematical reasoning, whereas certain inputs such as straightforward QA queries need only a small set of KV cache. Real-world inputs, however, are mixed and unpredictable, with no clear boundary by difficulty or domain.
Our propose objective is exactly to remove input-specific threshold constraints, motivating threshold-free methods that ensure consistent performance comparable to full-cache regardless of inputs, obviating the need of tuning for optimal thresholds. Pursuit of the best compression ratio is prioritized only after satisfying the objective. To the best of our knowledge, we are the first to propose an effective solution that fulfills such objective.
2.2 ReFreeKV
ReFreeKV consists of two stages implemented with efficient parallel operators. Conceptually, it first ranks all KV cache positions per layer and per attention head. It then sequentially retains key–value vectors until a stopping condition, determined by our input-insensitive metric with a universal threshold, is met; the KV cache at the remaining positions is subsequently discarded. In line with most prior KV compression works kv_survey, our method is applied once after input prefilling, with the primary goal of reducing memory consumption and an additional benefit of improved throughput (Appendix F).
Initial Ranking
The first stage ranks all KV cache, such that the beginning of the sequence may likely contain more critical information than its latter parts, which forms the basis for downstream sequential eviction.
Building on the positional bias reported in StreamingLLM DBLP:conf/iclr/XiaoTCHL24, we also observe that positions at the beginning and end of the input generally play a more critical role in subsequent generation. We exploit this property in the first stage, and rank KV cache by token positions as follows. Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head. The initial ranking takes the first $m$ positions and reversely takes the remaining nmn-m positions, denoted by $\widehat{X}=\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$. Note that $m$ is a chosen hyperparameter that works well regardless of specific input sequences.
Our early experiments also conduct exploration on other ranking strategies, reported in Section 3.4. By empirical validation, the position-based ranking is shown not only effective but also particularly advantageous in terms of computational overhead, constituting the first stage for ReFreeKV. However, pruning solely by positions does not fulfill our objective, as our experiments in Appendix G show that StreamingLLM can exhibit inconsistent performance varying budget sizes, highlighting the need for a more robust eviction strategy.
Eviction by Uni-Metric
With the initial ranking on KV cache, ReFreeKV then sequentially retains KV vectors, and halts upon the stopping condition by an input-insensitive Uni-Metric, after which the remaining cache is effectively evicted.
The design of Uni-Metric is then at the core of this process, which requires to signal the degradation level after removing KV cache of certain positions. we propose a metric that empirically correlates well with the performance change when discarding a position, which could serve as a bridge to ensure a minimal degradation upon the full cache performance. Inspired by DBLP:journals/corr/abs-2406-11430, we utilize the Frobenius norm (L2 norm) of the attention matrix $A\in\mathbb{R}^{n\times n}$ as the Uni-Metric, denoted as $A||_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|A_{i,j}|^{2}}$. For each position $i$ in the initially ranked sequence, we compare the Frobenius norm of the original attention matrix, $A||_{F}$, with that of a curated attention matrix, $\widetilde{A_{i}}||_{F}$, in which scores to all positions >i>i in the ranked sequence are masked out, replicating the effect of discarding all KV cache beyond position $i$. Once the norm difference reaches a threshold $T$ at position $i_{\text{prune}}$, the entire process terminates, retaining only the KV cache up to $i_{\text{prune}}$ and discarding the remainder, denoted as:
[…]
The Universal Threshold
To fulfill our objective, the threshold $T$ should ensure near lossless pruning invariant to inputs. Upon empirical search, we identify $T=1\%$ could serve well for this purpose. Figure 2 illustrates how performance varies with changes in the norm across different domains. Preliminary studies indicate that when $T<1\%$, performance remains comparable to the full-cache version robustly. We select $T=1\%$ to balance minimal degradation with maximal cache eviction. The efficacy of Uni-Metric and its universal threshold is validated at full scale in the main experiments presented in Section 3.2 and Figure 3.
Reducing Overhead